Creating and Using a Cost Formula

Big Thumbs Company manufactures portable flash drives for computers. Big Thumbs incurs monthly depreciation costs of $14,200 on its plant equipment. Also, each drive requires materials and manufacturing overhead resources. On average, the company uses 18,750 ounces of materials to manufacture 7,500 flash drives per month. Each ounce of material costs $3.00. In addition, manufacturing overhead resources are driven by machine hours. On average, the company incurs $30,000 of variable manufacturing overhead resources to produce 7,500 flash drives per month.

In your calculations, round variable rate per flash drive to the nearest cent. If required, round final answers to the nearest cent.

Required:

1. Create a formula for the monthly cost of flash drives for Big Thumbs.

Total cost of flash drives = Fixed cost  + ( Variable rate  x Number of flash drives)

Total cost of flash drives = $ + ($ x Number of flash drives)

2. If the department expects to manufacture 6,000 flash drives next month, what is the expected fixed cost (assume that 6,000 units is within the company's current relevant range)?
$

What is the total variable cost (assume that 6,000 units is within the company's current relevant range)?
$

What is the total manufacturing cost (i.e., both fixed and variable) (assume that 6,000 units is within the company's current relevant range)?
$
Answer #1

Total cost of flash drives = Fixed cost  + ( Variable rate  x Number of flash drives)

Fixed cost = 14,200

Variable cost = (18,750/7,500)*3 + (30,000/7,500)

= $11.5

Hence, Total cost of flash drives = $14,200 + ($11.5 x Number of flash drives)

2.Expected Fixed cost = $14,200

Note: Fixed costs do not change in the relevant range

total variable cost = 11.5*6,000 = $69,000

total manufacturing cost = Fixed cost + variable cost

= 14,200+69,000

= $83,200